An 80-pound grass seed mixture is 30% bluegrass seed. How many pounds of seed containing no bluegrass seed (0%) should be added to the seed mixture for a resulting blend that is 20% bluegrass seed? Construct a table, then solve
Answer by stanbon(75887)

An 80-pound grass seed mixture is 30% bluegrass seed. How many pounds of seed containing no bluegrass seed (0%) should be added to the seed mixture for a resulting blend that is 20% bluegrass seed?
------------------------------
Equation:
blue + blue = blue
0.30*80 + 0*x = 0.20(80+x)
Multiply thru by 100 to get:
30*80 = 20*80 + 20x
20x = 10*80
x = 40 lb. (amount of seed with no bluegrass that must be added to the mix)
===================================
Cheers,
Stan H.